### v1

# Vision-LSTM: xLSTM as Generic Vision Backbone

Abstract
Transformers are widely used as generic backbones in computer vision, despite initially introduced for natural language processing. Recently, the Long Short-Term Memory (LSTM) has been extended to a scalable and performant architecture – the xLSTM – which overcomes long-standing LSTM limitations via exponential gating and parallelizable matrix memory structure. In this report, we introduce Vision-LSTM (ViL), an adaption of the xLSTM building blocks to computer vision. ViL comprises a stack of xLSTM blocks where odd blocks process the sequence of patch tokens from top to bottom while even blocks go from bottom to top. Experiments show that ViL holds promise to be further deployed as new generic backbone for computer vision architectures.
Project page: https://nx-ai.github.io/vision-lstm/
1 Introduction
Language modeling architectures — such as Transformers [39, 1, 33] or more recently State Space Models [16, 17] such as Mamba [15] — are commonly adapted to the domain of computer vision to make use of their powerful modeling capabilities. However, in natural language processing, an input sentence is typically encoded into tokens that represent words or common subwords [7] via a discrete vocabulary. To encode images into a set of tokens, Vision Transformer [13] (ViT) proposed to group an input image into non-overlapping patches (of e.g. 16x16 pixel), linearly project them into a sequence of so-called patch tokens and add positional information to these tokens. This sequence can then be processed by language modeling architectures.
The Extended Long Short-Term Memory (xLSTM) family [4] was recently introduced as a new architecture for language modeling. It demonstrates the resurgence of LSTM in the LLM era, performing favorably against the likes of Transformers and State Space Models (SSMs). Analogous to existing vision versions of Transformers or State Space Models, e.g., ViT [13] or Vision Mamba [44], which have produced great results in various computer vision tasks [31, 22, 28, 30, 3], we introduce Vision LSTM (ViL) – a generic computer vision backbone that uses xLSTM blocks as its core components. To adjust xLSTM (an autoregressive model) to computer vision (an often non-autoregressive domain), we employ a stack of alternating mLSTM blocks [4] where odd blocks process patches row-wise from top left to bottom right and even blocks go from bottom right to top left. This simple alternating design allows ViL to efficiently process non-sequential inputs, such as images, without introducing additional computations.
Similar to vision adaptions of SSMs [23, 44], ViL exhibits linear computational and memory complexity w.r.t. sequence length which makes it appealing for tasks that benefit from high-resolution images such as medical imaging [8, 18, 38, 41], segmentation [22, 9], or physics simulations [5, 27, 6, 2]. In contrast, ViT’s computational complexity scales quadratically due to the self-attention mechanism, rendering them costly to apply to high-resolution tasks.
2 Method
Vision-LSTM (ViL) is a generic backbone for computer vision tasks, which is residually built from xLSTM blocks, as visualized in Figure 1. Following ViT [13], ViL first splits an image into non-overlapping patches via a shared linear projection, then adds learnable positional embeddings to each patch token. At the core of ViL are alternating mLSTM blocks, which are fully parallelizable and equipped with a matrix memory and a covariance update rule. Odd mLSTM blocks process patch tokens from top left to bottom right while even blocks go from bottom right to top left.
Figure 1: Schematic overview of Vision-LSTM (ViL). Following ViT [13], an input image is split into patches and linearly projected. Then, a learnable vector is added per position to the patches, producing a sequence of patch tokens. This sequence is then processed by alternating mLSTM blocks where even blocks flip the sequence before and after the mLSTM layer. For classification, ViL averages the first and the last patch followed by a linear classification head.
3 Experiments
We conduct experiments on ImageNet-1K [12], which contains 1.3M training images and 50K validation images where each image belongs to one of 1000 classes. Our comparison focuses on models that use a sequence modeling backbone and are roughly comparable in parameter counts.
We train ViL models for 800 epochs (tiny, tiny+) or 400 epochs (small, small+, base) on 224x224 resolution with a learning rate of 1e-3 using a cosine decay schedule. Detailed hyperparameters can be found in Appendix 5.
To facilitate a fair comparison against Vision Mamba (Vim) [44], we add additional blocks to our models to match parameter counts for the tiny and small variants (denoted as ViL-T+ and ViL-S+). Note that ViL requires significantly less compute than Vim since ViL traverses the sequence in an alternating way, whereas Vim traverses the sequence twice per block. This holds despite Vim using optimized CUDA kernels, which currently do not exist for mLSTM, and would speed-up ViL even more. We compare runtimes in Appendix A.1 where ViL is up to 69% faster than Vim.
As ViTs are well established in the vision community, they underwent multiple optimization cycles over the years [13, 34, 36, 35, 37, 19]. As this work is the first to apply xLSTM to computer vision, we do not expected to outperform years of hyperparameter tuning of ViTs in all cases. Even so, the results in Table 1 show that ViL shows favorable results to heavily optimized ViT protocols (DeiT, DeiT-II, DeiT-III) on smaller scales where only DeiT-III-S that was trained twice as long as ViL-S performs slightly better. On the “base” scale, ViL outperforms the initial ViT [13] models and achieves comparable results to DeiT [34]. Note that due to high costs of training models at this scale, the hyperparameters for ViL-B are far from optimal. For reference, training ViL-B takes roughly 600 A100 GPU-hours or 19 hours on 32 A100 GPUs.
The performance can be further improved by fine-tuning models in a “long-sequence fine-tuning” setting [44], which fine-tunes the models for 30 epochs with an increased sequence length of 729 by using 50% overlap between consecutive patch tokens.
ViL also demonstrates competitive performance with CNN-based models such as ConvNeXt [24], despite not leveraging the inductive bias inherent to convolutions.
Method #Params ACC
DeiT-T [34] 6M 72.2
DeiT-II-T [36] 6M 73.5
DeiT-III-T (reimpl.) 6M 76.2
Vim-T [44] 7M 76.1
ViL-T 6M 77.3
ViL-T+ 7M 78.1
DeiT-S [34] 22M 79.8
DeiT-II-S [36] 22M 80.7
DeiT-III-S [37] 22M 81.4
Vim-S [44] 26M 80.5
ViL-S 23M 80.7
ViL-S+ 26M 80.9
ViT-B384 [13] 86M 77.9
DeiT-B [34] 86M 81.8
DeiT-II-B [36] 86M 82.7
DeiT-III-B [37] 86M 83.7
ViL-B 88M 81.6*
Method #Params ACC
Long-sequence fine-tuning
DeiT-III-T (reimpl.) 6M 78.3
Vim-T [44] 7M 78.3
ViL-T+ 7M 80.0
DeiT-S [34] 22M 81.4
DeiT-III-S [37] 22M 83.2
Vim-S [44] 26M 81.6
ViL-S+ 26M 82.2
DeiT-B [34] 86M 83.2
DeiT-III-B [37] 86M 84.5
ViL-B 88M 82.7*
CNN-based models
ConvNeXt-T [24] 30M 82.2
ConvNeXt-S [24] 50M 83.1
ConvNeXt-B [24] 89M 83.8
SSM-based models
S4ND-Conv-T [26] 30M 82.2
S4ND-ViT-B [26] 89M 80.4
Table 1: ImageNet-1K classification accuracy. Entries in gray are trained trained for 400 epochs more than ViL. “Long-sequence fine-tuning” fine-tunes models with 50% overlapping patches for 30 epochs. We train DeiT models in the long-sequence fine-tuning setting as DeiT-II has no public checkpoints and DeiT-III only published checkpoints with longer training durations. *due to the large costs of training ViL-B, hyperparameters are far from optimal.
Block Design.
We study different ways to design ViL blocks in Table 2. The plain, uni-directional xLSTM block fails to reach competitive performances since the autoregressive nature of xLSTM is not suited for image classification. Traversing blocks in a bi-directional manner, i.e., introducing a second mLSTM layer in each block that traverses the sequence backwards (akin to Vim [44]), improves performance, but also requires more parameters and FLOPS. Sharing the parameters of the forward and backward mLSTM makes the model more parameter efficient, but still requires more compute and overloads these parameters which leads to performance drops. Using alternating blocks enhances performance while remaining compute and parameter efficient. We also explore quad-directional designs (similar to [23]), which refers to traversing the sequence row-wise (in both directions) and additionally column-wise (in both directions). Bi-directional traverses the sequence only row-wise (in both directions). Figure 2 visualizes the different traversal paths.
Figure 2: Visualization of traversal paths. Bi-directional uses the first two traversal paths. Quad-directional uses all of the above. Squares represent individual patch tokens.
Due to the increased cost of bi-directional and quad-directional blocks, this study was conducted in a heavily reduced setting. We train on a subset of ImageNet-1K that contains only samples from 100 classes for 400 epochs at 128x128 resolution. This was particularly necessary because our quad-directional implementation is not compatible with torch.compile (a generic speed optimization method from PyTorch [29]), which leads to longer runtimes, as indicated last column of Table 2. Due to this technical limitation, we choose the alternating bi-directional block as our core design.
Direction Alternating Param. Sharing ACC #Param FLOPS Runtime
Uni-dir. (xLSTM) 72.8 6.1M 1.0x 1.0x
Bi-dir. 74.8 6.5M 1.4x 1.9x
Bi-dir. 73.3 6.1M 1.4x 1.9x
Bi-dir. (ViL) 74.8 6.1M 1.0x 1.0x
Quad-dir. 76.3 7.3M 2.3x 7.5x
Quad-dir. 72.8 6.1M 2.3x 7.5x
Quad-dir. 75.6 6.1M 1.0x 1.9x
Table 2: Traversing the sequence of patch tokens in multiple directions yields better performance than the plain, uni-directional xLSTM block. However, using multiple directions per block also requires more compute and parameters. Sharing parameters between directions matches #Params but overloads them, leading to performance drops. Using multi-directional designs but alternating between them in consecutive blocks, matches #Params and FLOPS while achieving good performances. While quad-directional designs yields the best performances, they comes with drastically increased runtime. As our current quad-directional implementation is not compatible with torch.compile, it suffers from increased runtime even when alternating between directions. Since quad-directional designs are compute intensive, this study was conducted by training a ViL-T on an ImageNet-1K subset of 100 classes using resolution 128.
3.1 Classification Design
In order to perform classification with a ViT, the sequence of tokens is pooled into a single token, which is then used as input to a classification head. The most common methods to do this pooling are (i) adding a learnable [CLS] token at the start of the sequence or (ii) averaging all patch tokens to produce an [AVG] token. Whether to use the [CLS] or [AVG] token is typically a hyperparameter, where both variants achieve comparable performances. On the contrary, autoregressive models often require specialized classification designs. For example, Vim [44] requires the [CLS] token to be in the middle of the sequence, suffering heavy performance losses if other classification designs, e.g., an [AVG] token or two [CLS] tokens at start and end of the sequence, are employed. Due to its autoregressive nature, we explore different classification designs for ViL in Table 3. [AVG] averages all patch tokens, “Middle Patch” uses the middle patch token, “Middle [CLS]” uses a [CLS] token in the middle of the sequence, “Bilateral [AVG]” uses the average of the first and last patch token. We find ViL to be relatively robust the classification design where all performances are within 0.6%. We choose “Bilateral [AVG]” over “Middle [CLS]”, as ImageNet-1K is known to have a center bias, i.e., objects are commonly located in the middle of the picture. By using “Bilateral [AVG]” we avoid exploiting this bias, which makes our model more general.
Arch ACC
[AVG] 75.4
Middle Patch 75.9
Middle [CLS] 75.3
Bilateral [AVG] 75.9
Bilateral Concat 76.4
Table 3: How to pool the sequence of patch tokens into a single token for classification? ViL is relatively robust to the classification design. Using the middle patch token and averaging the first and last token (“Bilateral [AVG]”) work similarly well. The good performance of middle patch could be caused by the bias in ImageNet-1K that objects are typically in the center of the image. We therefore choose “Bilateral [AVG]” to avoid exploiting this bias. “Bilateral Concat” concatenates the first and last patch token which introduces additional parameters in the classification head. To save compute, we conduct this study by training a ViL-T on ImageNet-1K for 400 epochs in 192x192 resolution followed by fine-tuning for 20 epochs on 224x224 resolution.
To remain comparable to previous architectures that use a single token as input to the classification head, we average the first and the last patch. To achieve best performances, we suggest to concatenate the two tokens (“Bilateral Concat”) instead of averaging them. This is similar to common practices in self-supervised vision transformers, such as DINOv2 [28], which are trained with two objectives that are attached at the [CLS] and [AVG] token respectively and therefore benefit from concatenating the representation of [CLS] and [AVG] token. This directions has also been explored for vision SSM models [40] where multiple [CLS] tokens are spread out over the sequence and then used as input to a classifier. Similar directions could also improve the performance of ViL.
4 Conclusion
Motivated by the success of xLSTM in language modeling, we introduced ViL, an adaption of the xLSTM architecture to vision tasks. ViL processes a sequence of patch tokens in alternating fashion. Odd blocks process image patches row-wise from top left to bottom right and even blocks go from bottom right to top left. Our new architecture outperforms SSM-based vision architectures and also optimized ViT models on ImageNet-1K classification. Remarkably, ViL is able to outperform ViT training pipelines, that resulted from years of hyperparameter tuning and transformer improvements, in fair comparisons.
In the future, we see potential in applying ViL when high-resolution images are needed for optimal performance, such as semantic segmentation or medical imaging. In these settings, transformers suffer from high computational costs due to the quadratic complexity of self-attention, which ViL does not have due to its linear complexity. Additionally, improving pre-training schemes (e.g. via self-supervised learning), exploring better hyperparameter settings or transferring techniques from transformers (e.g. LayerScale [35]) are promising directions of ViL.
Acknowledgments
We acknowledge EuroHPC Joint Undertaking for awarding us access to Karolina at IT4Innovations, Czech Republic, MeluXina at LuxProvide, Luxembourg, Leonardo at CINECA, Italy and LUMI at CSC, Finland.
The ELLIS Unit Linz, the LIT AI Lab, the Institute for Machine Learning, are supported by the Federal State Upper Austria. We thank the projects Medical Cognitive Computing Center (MC3), INCONTROL-RL (FFG-881064), PRIMAL (FFG-873979), S3AI (FFG-872172), DL for GranularFlow (FFG-871302), EPILEPSIA (FFG-892171), AIRI FG 9-N (FWF-36284, FWF-36235), AI4GreenHeatingGrids (FFG- 899943), INTEGRATE (FFG-892418), ELISE (H2020-ICT-2019-3 ID: 951847), Stars4Waters (HORIZON-CL6-2021-CLIMATE-01-01). We thank Audi.JKU Deep Learning Center, TGW LOGISTICS GROUP GMBH, Silicon Austria Labs (SAL), FILL Gesellschaft mbH, Anyline GmbH, Google, ZF Friedrichshafen AG, Robert Bosch GmbH, UCB Biopharma SRL, Merck Healthcare KGaA, Verbund AG, GLS (Univ. Waterloo), Software Competence Center Hagenberg GmbH, Borealis AG, TÜV Austria, Frauscher Sensonic, TRUMPF and the NVIDIA Corporation.
Appendix A Extended Results
A.1 Runtime Comparison of ViL vs Vim
We compare the time it takes to train ViL and Vim [44] for a single ImageNet-1K epoch in Table 4. We follow the scaling procedure of ViTs, using 192 (T), 384 (S), 768 (B), 1024 (L) as latent dimension where the large scale additionally doubles the number of blocks.
Optimization (T)iny (S)mall (B)ase (L)arge
Vim [44] custom CUDA kernel 7.3h 14.0h 28.2h 76.4h
ViL torch.compile 5.0h 8.7h 16.6h 45.1h
Speedup of ViL compared to Vim 45% 61% 69% 69%
Table 4: Runtime comparisons between Vim [44] and ViL. ViL is up to 69% faster despite the current lack of a custom CUDA kernel. As mLSTM (and ViL) can be parallelized analogous to FlashAttention [11, 10] via custom hardware optimizations, ViL will become even faster in the future. Runtimes denote the training time for 10 ImageNet-1K epochs and are extrapolated from short benchmark runs on a single A100-80GB-PCIe using float16 precision and 224x224 images.
A.2 Impact of Longer Training
We investigate training for a longer duration on ViL-T+ and train it for 400 and 800 epochs following the hyperparameters in Table 5. After 400 epochs, ViL-T+ achieves 77.2% accuracy and after doubling the training duration the model achieves 78.1% accuracy.
Our implementation of DeiT-III-T achieves 75.6% after 400 epochs and 76.2% with 800 epochs using the hyperparameters in Table 6.
Appendix B Implementation Details
B.1 Hardware
We train models on a mix of custom hardware servers (mainly A100 and A40 GPUs) and public research clusters equiped with either 8xA100 or 4xA100 nodes.
We estimate the total number of A100 GPU-hours used for this project to be 25K hours. This estimate includes everything from initial exploration, method development, analysis and evaluations.
B.2 FLOPS Calculation
We calculate the FLOPS from Table 2 with the fvcore library [^1] https://github.com/facebookresearch/fvcore. As fvcore does not support all operations of ViL, the FLOPS are not 100% accurate, but nevertheless serve as a good reverence point to compare the relative computational effort between block designs.
B.3 ViL Hyperparameters
Table 5 shows detailed hyperparameters used to train ViL models. In the “long-sequence fine-tuning” setting, we follow Vim [44] and fine-tune for 30 epochs with a learning rate of 1e-5, a weight decay of 1e-8 using an overlap of 50% (8 pixels) between consecutive patches. This increases the sequence length from 196 to 729. Other hyperparameters follow Table 5.
Parameter Value
Epochs 800 (T), 400 (S/B)
Batch size 1024 (S), 2048 (T/B)
Model
   Patch size 16x16
   Latent dimension 192 (T), 384 (S), 768 (B)
   Depth 24 (T, S, B), 29 (T+), 26 (S+)
Stochastic depth [20]
   Peak rate 0 (T), 0.05 (S), 0.2 (B)
   Layer-wise Decay
Optimizer AdamW [25, 21]
   Base Learning rate 1e-3
   Linear LR Scaling Divisor [14] 1024
   Weight decay 0.05
   Momentum $\beta_{1}=0.9,\beta_{2}=0.999$
   Gradient Norm Clip 1.0
Learning rate schedule linear warmup $\rightarrow$ cosine decay
   Warmup epochs 5
   End LR 1e-6
Label smoothing [32]
Train Data Augmentation
   RandomResizedCrop 224
    Scale [0.08, 1.0]
    Interpolation bicubic
   RandomHorizontalFlip $p=0.5$
   3-Augment [37]
    Gaussian Blur $\sigma$ [0.1, 2.0]
    ColorJitter [0.3, 0.3, 0.3, 0.0]
   Normalize ImageNet-1K statistics
   Mixup [43] $\alpha$ 0.8
   Cutmix [42] $\alpha$ 1.0
Test Data Augmentation
   Resize 224
    Interpolation bicubic
   CenterCrop 224
   Normalize ImageNet-1K statistics
Table 5: Hyperparameters for training ViL on ImageNet-1K, inspired by DeiT-III [37].
B.4 DeiT-III Reimplementation Hyperparameters
Table 5 shows detailed hyperparameters used to train DeiT-III-T (reimpl.) from Table 1.
Parameter Value
Epochs 800 $\rightarrow$ 20
Batch size 2048
Stochastic depth [20]
Layerscale [35] 1e-4
Optimizer AdamW [25, 21]
   Base Learning rate 1e-3
   Linear LR Scaling Divisor [14] 1024
   Weight decay 0.05
   Momentum $\beta_{1}=0.9,\beta_{2}=0.999$
   Gradient Norm Clip
Learning rate schedule linear warmup $\rightarrow$ cosine decay
   Warmup epochs 5
   End LR 1e-6
Label smoothing [32]
Train Data Augmentation
   RandomResizedCrop 192 $\rightarrow$ 224
    Scale [0.08, 1.0]
    Interpolation bicubic
   RandomHorizontalFlip $p=0.5$
   3-Augment [37]
    Gaussian Blur $\sigma$ [0.1, 2.0]
    ColorJitter [0.3, 0.3, 0.3, 0.0]
   Normalize ImageNet-1K statistics
   Mixup [43] $\alpha$ 0.8
   Cutmix [42] $\alpha$ 1.0
Test Data Augmentation
   Resize 192 $\rightarrow$ 224
    Interpolation bicubic
   CenterCrop 192 $\rightarrow$ 224
   Normalize ImageNet-1K statistics
Table 6: Hyperparameters for training our reimplementation of DeiT-III-T [37] on ImageNet-1K. The most significant change is that we reduce the learning rate from 3e-3 to 1e-3 as we found this to greatly improve performance. We make minor changes to the protocol such as using AdamW or no gradient clipping as models were stable without it. We follow the best setting from DeiT-III [37] and pre-train on 192 resolution followed by a short fine-tuning on 224 resolution (indicated by $\rightarrow$).